### v1

# Motion-Aware Generative Frame Interpolation

Abstract
Generative frame interpolation, empowered by large-scale pre-trained video generation models, has demonstrated remarkable advantages in complex scenes. However, existing methods heavily rely on the generative model to independently infer the correspondences between input frames—an ability that is inadequately developed during pre-training. In this work, we propose a novel framework, termed Motion-aware Generative frame interpolation (MoG), to significantly enhance the model’s motion awareness by integrating explicit motion guidance. Specifically we investigate two key questions: what can serve as an effective motion guidance, and how we can seamlessly embed this guidance into the generative model. For the first question, we reveal that the intermediate flow from flow-based interpolation models could efficiently provide task-oriented motion guidance. Regarding the second, we first obtain guidance-based representations of intermediate frames by warping input frames’ representations using guidance, and then integrate them into the model at both latent and feature levels. To demonstrate the versatility of our method, we train MoG on both real-world and animation datasets. Comprehensive evaluations show that our MoG significantly outperforms the existing methods in both domains, achieving superior video quality and improved fidelity.
Figure 1: Examples of frame interpolation in real-world and animation scenes. Compared to other methods, our approach, MoG, exhibits superior stability in motion and consistency in content relative to the input frames. For optimal viewing, please zoom in.
*Work is done during internship at Tencent PCG.  ‡Equal contribution.  †Corresponding author (lmwang@nju.edu.cn).
1 Introduction
Video Frame Interpolation (VFI), which seeks to synthesize intermediate frames between two input frames, has garnered significant attention in recent years due to its capacity for enhancing video frame rates in video post-processing. Flow-based VFI methods [10, 41, 42, 15, 13] predominantly rely on correspondence information through estimating the motion between input frames—termed intermediate flow—to warp information and generate intermediate frames. However, these approaches are primarily optimized for rigid motion scenarios. Consequently, when confronted with complex scenes, the interpolated frames often exhibit pronounced blurring and artifacts, as the results of the flow-based method EMA-VFI [41] shown in Fig. 1.
To overcome these limitations, recent advancements [34, 35, 33, 3] have shifted towards leveraging video generation models [35, 2] for frame interpolation, capitalizing on their strong generative capabilities in dynamic scenes. Despite the notable improvements exhibited in complex scenarios [34], current approaches still face a key challenge: they fail to explicitly exploit the dynamics between the two input frames. Instead, they solely rely on the generative model to infer the correspondences between input frames by itself—an ability that is insufficiently nurtured during the generative pre-training. Consequently, the videos generated by these methods often suffer from unstable motion and incoherent content to input frames, as shown by DynamiCrafter [35] and ToonCrafter [34] in Fig. 1.
In this work, we introduce a new framework, Motion-aware Generative frame interpolation (MoG), marking the first explicit incorporation of motion guidance between input frames to enhance the motion awareness of generative models. Our approach significantly alleviates the difficulty of inferring the dynamics between input frames to generate the realistic motion. To achieve this goal, we address two pivotal questions: what can serve as an effective motion guidance, and how to seamlessly and harmlessly integrate this guidance into the generative model.
For the first question, we propose leveraging the intermediate flow from flow-based frame interpolation models [41] as explicit motion guidance. The intermediate flow can be utilized to obtain coarse approximations of intermediate frames under the assumption of rigid motion, which can serve as effective guidance to inspire and stabilize the inference of realistic motion in the generation process. Moreover, compared to direct estimation by pre-trained optical flow models [36, 28, 37], the intermediate flow is task-oriented trained [38], making it inherently more suitable for generating intermediate frames. Flow-based interpolation models also specifically account for occlusion masks [41] between the two input frames, enabling a more accurate aggregation of information into the intermediate frame—something that optical flow often fails to achieve.
To address the second question, we propose a simple yet effective strategy: warping the information from input frames to derive the guidance-based intermediate frame representations, which are subsequently integrated into the pre-trained model as explicit motion cues. This integration is performed at both the latent and feature levels, enabling the model to be fully aware of the correspondences between input frames. Specifically, at the latent level, we concatenate a collection of auxiliary latents to the input at each denoising step [8], which includes the latents from the input frames as well as the warped intermediate latents. In terms of feature-level integration, we reuse the pre-trained temporal layers [35] to temporally smooth the warped features, ensuring their alignment within the original feature space, and then incorporate them into the network. Notably, our design requires no additional parameters and achieves improved performance with few fine-tuning steps.
To comprehensively evaluate the versatility of our method, we develop MoG for both the real-world and animation scenes. Furthermore, we meticulously curate two testing datasets, VFIBench-Real and VFIBench-Ani, to assess the performance of each scenes. Experimental results indicate that MoG significantly outperforms the existing generative interpolation models in both qualitative and quantitative aspects. Qualitatively, the videos generated by MoG exhibit enhanced motion stability and greater content consistency, as evidenced in Fig. 1. Quantitatively, MoG demonstrates the obvious improvements in video quality metrics and superior fidelity when compared to ground truth videos. Our contributions are summarized as follows:
  • We introduce a new generative frame interpolation framework, MoG, which is the first to incorporate motion guidance to enhance the motion awareness.
  • We demonstrate that the intermediate flow from flow-based VFI can effectively serve as a motion guidance, and we design a simple and efficient method to integrate this prior into the network at both the latent and feature levels.
  • We validate the effectiveness of MoG in both real-world and animated scenarios, with experimental results showing that MoG significantly outperforms existing open-source generative frame interpolation methods.
2 Related Work
Figure 2: Overview of MoG. MoG consists of two parts. First, it extracts the motion guidance of the input frames using a pre-trained flow-based VFI model. Subsequently, this guidance is seamlessly injected into the generative model at both the latent and feature levels.
2.1 Flow-based Frame Interpolation
Flow-based video interpolation, explicitly estimating the intermediate flow from the input frames to the intermediate frame, have become dominant in deterministic frame interpolation [17]. It can be broadly categorized into two classes based on how the intermediate optical flow is derived. The first class [9, 19, 18, 20, 1, 12] utilized pre-trained optical flow models to obtain the intermediate flow either directly or through refinement. For instance, SoftSplat [19] linearly adjusted the bidirectional flow estimated by PWC-Net [27] to represent the intermediate flow and employs an improved forward warping to aggregate information. The second class [41, 10, 42, 15, 13, 14, 21, 16] modeled the correspondence information of the input frames to directly predict the intermediate flow, offering greater flexibility and task-oriented modeling capacity compared to the first approach. RIFE [10] demonstrated that simple convolutional layers can effectively predict the intermediate flow, achieving impressive efficiency. Similarly, EMA-VFI [41] enhanced the flow prediction by explicitly modeling the dynamics between frames through inter-frame cross-attention. However, both of classes are primarily designed under the assumption of rigid motion scenes. When confronted with complex motion scenarios, they often exhibit significant blurring and artifacts. In our work, we focus on the generative frame interpolation by leveraging the second class of flow-based video frame interpolation (EMA-VFI [41]) to provide explicit motion guidance.
2.2 Generative Frame Interpolation
Recent work has begun to explore the use of large-scale pre-trained video generation models [2, 35], which excel at generating videos in complex dynamic scenes, for the VFI task. Current generative frame interpolation methods can be categorized into two types: the first [3, 33] employed pre-trained generative models to perform image-to-video tasks conditioned on the initial and final frames, subsequently merging the resulting videos to create the final interpolated frame. For example, GI [33] enhanced the motion stability by controlling the consistency of temporal correlations across the two generation processes. The second category [35, 34, 15, 32, 44] focused on fine-tuning video generation models specifically for interpolation, by integrating information from input frames into the model’s architecture and optimizing it for end-to-end interpolation. DynamiCrafter [35] was trained for real-world interpolation, while ToonCrafter [34] was tailored for animated scenes. Although all these methods have demonstrated significant improvements in generating complex scenarios, they do not explicitly consider the correspondence between input frames, which complicates the motion inference of generative models. In contrast, we are the first to explicitly introduce the correspondence guidance to enhance the motion awareness of generative models and our method achieves superior video quality and fidelity.
3 Preliminaries
Figure 3: Comparison of different methods for approximating intermediate frames. Linear interpolation merely overlays the input frames, while optical flow often introduces significant artifacts. In contrast, the frames warped through intermediate flow demonstrate superior stability, providing the generative model with robust correspondence information.
3.1 Task Definition
For the input frames $x^{1}$ and $x^{N}\in\mathbb{R}^{3\times H\times W}$, frame interpolation aims to generate a video comprising $N$ frames, denoted as $\mathbf{x}\in\mathbb{R}^{N\times 3\times H\times W}$, where the first and last frames correspond to the input frames.
3.2 Intermediate Flow from Flow-based VFI
Flow-based methods explicitly estimate the correspondence between the starting and ending frames with respect to the intermediate frame, termed the intermediate flow. The intermediate flow can be obtained either by scaling the optical flow between frames [9, 19] or through direct prediction [10, 41]. In this work, we adopt the prediction-based method EMA-VFI [41], owing to its versatility across various time steps and its task-oriented training [39].
Specifically, given the input frames $x^{1},x^{N}\in\mathbb{R}^{3\times H\times W}$ as well as the $n$-th frame $x^{n}$ to be predicted, the intermediate flow $f$ is computed using a learnable network $\mathcal{O}$:
$f_{1\to n},f_{N\to n},M^{n}=\mathcal{O}(x^{1},x^{N},n).$ (1)
Here, $f_{i\to n}\in\mathbb{R}^{2\times H\times W}$ denotes the intermediate flow from the $i$-th frame $x^{i}$ to the $n$-th frame $x^{n}$, and $M\in\mathbb{R}^{1\times H\times W}$ represents the occlusion mask between the two frames at the $n$-th frame, taking values in the range of $(0,1)$. Subsequently, we can coarsely estimate the intermediate frame $\bar{x}^{n}$ as follows:
$\bar{x}^{n}=\text{warp}(x^{1},f_{1\to n})\odot M^{n}+\text{warp}(x^{N},f_{N\to n})\odot(1-M^{n}),$ (2)
where $\text{warp}(x^{i},f_{i\to n})$ denotes the backward warping by $f_{i\to n}$, and $\odot$ signifies the element-wise multiplication.
3.3 VFI with Diffusion Models
Empowered by large-scale pre-training, video diffusion models [2, 35] exhibit remarkable capabilities in generating videos within complex scenarios. Recent works [33, 34, 35, 3] have begun to leverage pre-trained video diffusion models for frame interpolation tasks. In this work, we explore our method based on two generative frame interpolation models, DynamiCrafter [35] and ToonCrafter [34], for real-world and animation scenes respectively. Both models are based on Latent Diffusion Models (LDMs) [25], which conduct diffusion in the latent space of an autoencoder. Specifically, for any video $\mathbf{x}\in\mathbb{R}^{N\times 3\times H\times W}$, where $N$ denotes the number of frames, the video is transformed into the latent space using a pre-trained encoder $\mathcal{E}$ (i.e., VQ-VAE) [25] to obtain the corresponding latent code $\mathbf{z}_{0}=\mathcal{E}(\mathbf{x})\in\mathbb{R}^{N\times C\times h\times w}$.
During training, $\mathbf{z}_{0}$ is first converted into an intermediate noisy video at timestep $t$ using the equation:
$\mathbf{z}_{t}=\alpha_{t}\mathbf{z}_{0}+\sqrt{1-\alpha_{t}}\mathbf{\epsilon},\quad\mathbf{\epsilon}\sim\mathcal{N}(0,I).$ (3)
To achieve frame interpolation task, a learnable denoising network $\epsilon_{\theta}$ is then employed to predict the noise $\epsilon$ given the condition information from the first and the last frames. DynamiCrafter and ToonCrafter incorporate such condition information by:
$\mathbf{\tilde{z}}_{t}=\left[\mathbf{z_{t}};\mathbf{\bar{z}}_{0}\right],\quad\mathbf{\tilde{z}}_{t}\in\mathbb{R}^{N\times(2\times C)\times h\times w},$ (4)
where $\mathbf{\bar{z}}_{0}$ is composed of the latent codes of bound frames $z_{0}^{1},z_{0}^{N}$, while other positions remain zero. Then the denoising network is optimized by minimizing the following loss:
$\mathcal{L}(\theta)=\mathbb{E}_{\mathbf{z}_{0},t,\epsilon\sim\mathcal{N}(0,I)}\left[\left\|\epsilon-\epsilon_{\theta}\left(\mathbf{\tilde{z}}_{t};t,c\right)\right\|_{2}^{2}\right].$ (5)
Here, $c$ includes other condition information like the text and the fps. After training, we can iteratively recover $\mathbf{\hat{z}}_{0}$ using the input conditions and pure noise $\mathbf{z}_{T}\sim\mathcal{N}(0,I)$, generating the video $\mathbf{\hat{x}}=\mathcal{D}(\mathbf{\hat{z}}_{0})$ via the decoder $\mathcal{D}$.
The design of the denoising network $\epsilon_{\theta}$ follows an U-Net-like structure [26], consisting of contracting blocks, middle blocks, and expansive blocks. Each block comprises spatial and temporal layers. The spatial layers mainly consist of ResNet blocks [6] and Transformer blocks [31] with spatial attention, modeling spatial information within each frame, while the temporal layers are formed by Transformer blocks with temporal self-attention.
4 Motion-Aware Generative VFI
Currently, most generative interpolation methods [33, 3, 35, 34, 5] conduct frame interpolation by directly incorporating information from the input frames into the model’s input [5, 35] or decoder [35]. However, in these methods, the correspondence between input frames, which is crucial for understanding and reasoning about the motion and appearance of the intermediate frames, have to be inferred independently by the generative model itself. This correspondence estimation capability has not been adequately cultivated in the pre-training of large-scale text-to-video generation or pure video generation tasks. Consequently, this insufficient exploration of correspondence leads to inconsistencies in motion and appearance of the generated videos relative to the input frames.
To address this issue, we propose Motion-Aware Generative Frame Interpolation (MoG), the first approach to explicitly introduce correspondence information between the two frames into the generative model, thereby reducing the difficulty of the inference for motion and appearance for generative model. As illustrated in Fig. 2, MoG comprises two parts: it first extracts an explicit motion guidance based on the input frames by a pre-trained flow-based VFI model [41]; it then seamlessly injects this guidance into the generative model at both the latent and feature levels. Notably, our method does not require any additional parameters; it merely necessitates minimal fine-tuning steps on spatial layers of the denoising network to compensate for changes in feature distribution.
4.1 Choice of motion guidance
As illustrated in Fig. 3, in addition to the intermediate flow described in Eq. 1, there are two other guidance for representing intermediate frames based on input frames. The first method is linear interpolation, which directly weights the representations of the initial and final frames according to the time step of the intermediate frame. While straightforward, this technique merely overlays the input frames, offering limited correspondence information. The second method employs a pre-trained optical flow estimator, such as [37], to derive the flow between input frames, which is then scaled to generate the intermediate flow. This approach often suffers from limited generalization due to being predominantly trained on synthetic datasets. Moreover, it cannot account for occlusion relationships between input frames. This can lead to artifacts in complex scenes, such as jagged edges and ghosting effects.
In contrast, the intermediate flow defined in Eq. 1 is specifically trained for the VFI task, making it inherently more suitable for representing the correlation between intermediate frames and the initial and final frames. It also predicts an occlusion mask $M^{n}$ for each frame. As illustrated in Fig. 3, the intermediate flow allows for the approximation of intermediate frames, which, despite some blurriness, maintains stable motion trajectories and consistent object appearances. This stability alleviates the difficulties of real motion inference in generative models. This is further validated by our experiments in Sec. 5.4.
4.2 Motion Guidance Integration
Due to the inability of diffusion models to directly use correspondence guidance, it is necessary to devise a strategy to inject these motion informaiton into the denoising network. To achieve this, we propose a simple yet effective approach that leverages the guidance, akin to the Eq. 2, to coarsely estimate the representation of the intermediate frames from the representations of the start and end frames. This estimated representation is then seamlessly merged into the denoising network. Furthermore, to fully exploit the correspondence guidance across different granularities of the generation process, we conduct the guidance injection at both the latent and feature levels.
Latent-Level Injection.
To introduce motion guidance at the latent level, we propose to coarsely estimate the latent code of intermediate frames by backward warping of the input frame’s latent code with the intermediate flow. Specifically, given the latent codes of the start and end frames $z_{0}^{1}$ and $z_{0}^{N}$, along with the motion guidance obtained via Eq. 1, we estimate the latent code of the $n$-th intermediate frame:
$\bar{z}_{0}^{n}=\text{warp}(z_{0}^{1},f_{1\to n})\odot M^{n}+\text{warp}(z_{0}^{N},f_{N\to n})\odot(1-M^{n}).$ (6)
Here, $\bar{z}_{0}^{n}$ represents the estimated latent code for the $n$-th frame using the motion guidance. During training, the input to the denoising network is modified to:
$\tilde{\mathbf{z}}_{t}=\left[\mathbf{z}_{t};\bar{\mathbf{z}}_{0}\right],\quad\tilde{\mathbf{z}}_{t}\in\mathbb{R}^{N\times(2\times C)\times h\times w}.$ (7)
It is noteworthy that no additional parameters are required, as our base models, DynamiCrafter or ToonCrafter, have already designed to accommodate extra inputs; however, the latent codes of intermediate frames in $\mathbf{\bar{z}}_{0}$ are always zero at their methods.
Feature-Level Injection.
To effectively integrate the motion guidance in different granularities, we propose to inject guidance also in feature-level. Analogous to the latent-level, we estimate the features of intermediate frames based on the features $F^{0},F^{N}\in\mathbb{R}^{D\times H\times W}$ of the input frames:
$\bar{F}^{i}=\text{warp}(F^{1},f_{1\to i})\odot M^{n}+\text{warp}(F^{N},f_{N\to i})\odot(1-M^{n}).$ (8)
In this equation, $\bar{F}^{i}$ represents the estimated features of the $i$-th frame under the correspondence prior. Unfortunately, unlike in latent-level injection, direct concatenation of the warped intermediate features into the network is not feasible. To address this issue, we leverage the temporal layer of the denoising network, which enables us to smooth and align the estimated intermediate features with the original feature distribution without introducing additional parameters, as shown in Fig. 2:
[…]
Subsequently, we incorporate the smoothed features $\dot{F}$ into the original features:
$\mathbf{\tilde{F}}=(\mathbf{F}+\mathbf{\hat{F}})/2.$ (10)
Remarkably, our exploration (refer to Sec. 5.4) reveals that this simple averaging already allows the generative model to effectively utilize the introduced motion guidance.
5 Experiment
[…]
Average
Real Anime Real Anime Real Anime Real Anime Real Anime Real Anime Real Anime
      Flow-based VFI models
RIFE [10] 86.63 86.92 91.67 94.25 97.92 98.46 99.28 99.38 45.18 48.22 46.15 48.11 77.81 79.22
EMA-VFI [41] 87.85 88.97 91.91 95.12 97.39 98.29 99.11 99.29 46.68 50.20 47.96 52.24 78.48 80.69
      Generative VFI models
GI [33] 91.81 91.28 91.70 94.16 93.86 96.43 97.51 98.00 47.44 51.13 57.51 60.84 79.97 81.97
TRF [3] 89.25 89.25 92.26 94.01 93.48 95.77 96.77 97.18 47.94 50.12 54.38 58.40 79.01 80.79
DynamiCrafter [35] 89.06 91.94 92.97 96.39 48.47 58.24 79.51
ToonCrafter [34] 91.78 95.42 95.63 96.75 51.44 64.37 82.57
MoG (ours) 92.65 92.91 94.34 96.11 94.69 96.58 97.90 96.94 49.84 52.43 59.23 64.89 81.44 83.31
Table 1: Quantitative comparison on video quality. Bold text indicates the best results in generative frame interpolation.
[…]
Real Anime Real Anime Real Anime Real Anime Real Anime Real Anime
      Flow-based VFI models
RIFE [10] 18.13 20.27 0.6094 0.7817 0.3745 0.3407 898.38 647.49 65.04 66.16 0.8228 0.8643
EMA-VFI [41] 18.11 20.45 0.6118 0.7843 0.3763 0.3701 808.54 544.48 60.28 55.84 0.8363 0.8895
      Generative VFI models
GI [33] 15.95 18.04 0.5271 0.6971 0.3384 0.2891 521.00 449.31 36.06 46.18 0.8703 0.8710
TRF [3] 15.43 16.49 0.5132 0.6744 0.3920 0.3470 624.63 481.02 42.48 53.95 0.8491 0.8731
DynamiCrafter [35] 16.05 0.5225 0.3380 562.34 42.16 0.8634
ToonCrafter [34] 18.01 0.7182 0.2944 425.71 40.63 0.9203
MoG (ours) 17.82 19.44 0.5898 0.7434 0.2716 0.2615 401.49 351.41 31.26 33.73 0.9083 0.9320
Table 2: Quantitative comparison on generation fidelity to the ground truth.
5.1 Implementation Details
We develop MoG based on DynamiCrafter [35] for real-world scenes and ToonCrafter [34] for animation scenes. MoG employs EMA-VFI [41] for intermediate flow prediction. For model fine-tuning, we only train the spatial layers, while keeping all other parameters fixed. We train with the same loss in Eq. 5 for 20K steps on $1\times 10^{-5}$ learning rate and batch size 32. The training dataset is internal collected of $512\times 320$ resolution with 16 frames. The sampling strategy is consistent with [35] and [34].
5.2 VFIBench
To evaluate interpolated frames, we present VFIBench, a comprehensive benchmark that encompasses diverse data, including real-world videos and animations. It employs various metrics for a detailed assessment of frame quality and fidelity to ground truth. VFIBench also poses a challenge by requiring models to interpolate 14 frames between specified start and end frames. This setup demands advanced motion modeling capabilities. For data collection, we meticulously selected 100 samples from the DAVIS 2017 dataset [23], referred to as the VFIBench-Real, to reflect real-world scenarios. Additionally, we curate another set of 100 samples from internet animations, called VFIBench-Ani, which includes a diverse range of styles from Japanese, American, and Chinese animations.
A well-interpolated video should not only be of high quality inherently but also maintain fidelity to the ground truth. For the former, we utilize six metrics from VBench [11]: subject consistency, background consistency, temporal flickering, motion smoothness, aesthetic quality, and imaging quality. These metrics collectively assess the intrinsic quality of the video. For the latter, we employ six widely adopted metrics: PSNR, SSIM, LPIPS [43], FID [7], and the CLIP similarity score [24] for image-level comparison, and FVD [29, 30] for video-level comparison.
Figure 4: Visual comparison on real-world and animation scenes.
5.3 Comparative Analysis
We evaluate our MoG by benchmarking it against state-of-the-art methods across two categories: flow-based interpolation methods, specifically RIFE [10] and EMA-VFI [41], and generative interpolation methods, including GI [33], TRF [3], DynamiCrafter [35] and ToonCrafter [34]. Note that DynamiCrafter and ToonCrafter are tailored for real-world and cartoon animation data, respectively, and their performance is reported separately for each data type.
Quantitative results.
As shown in Tab. 1 and Tab. 2, compared to others generative VFI methods, MoG exhibits significant improvements in video quality and fidelity to ground truth, particularly in consistency-related metrics. This demonstrates that the introduced motion guidance indeed enables the generative model to better understand the correspondence between input frames. Compared to flow-based VFI models, our approach also demonstrates notable enhancements across most metrics; however, it lags in PSNR, SSIM, Temporal Flickering, and Motion Smoothness. We argue that this discrepancy arises because flow-based VFI often produces blurry results in complex motion scenarios (as illustrated in Fig. 1), which can inflate these metrics while compromising actual visual quality [43].
Qualitative results.
We present qualitative comparisons with three generative VFI methods in Fig. 4. Lacking explicit motion guidance, these methods struggle to accurately infer and understanding the correspondences between input frames, resulting in inconsistent content and unstable motion. In contrast, MoG achieves superior motion and visual quality in complex scenarios. More comparisons are available on the provided website in supplementary materials.
User study.
To further verify the advantage of our method, we also conduct a comprehensive user study. Participants are instructed to select the best-generated videos based on motion quality, temporal coherence, frame fidelity, and overall quality. We collect results from 27 participants and report the findings in Tab. 3. Thanks to the explicit motion guidance, the study shows a clear preference for our method in all aspects.
5.4 Ablation Study
For brevity, we only conduct ablation experiments in real-world scenarios. Our analysis primarily relies on four metrics to evaluate different strategies: two pertaining to video quality, namely Subject Consistency and Background Consistency (abbreviated as Sub. Cons. and Bg. Cons. in Tab. 4), and two metrics assessing fidelity between the video and ground truth, specifically LPIPS and FVD.
[…]
Quality
  EMA-VFI [41] 0.49% 0.74% 0.49% 0.49%
TRF [3] 1.73% 1.73% 0.99% 1.48%
GI [33] 16.05% 14.57% 23.21% 15.56%
DynamiCrafter [35] 3.70% 3.21% 4.20% 2.22%
MoG (ours) 78.02% 79.75% 71.11% 80.25%
Table 3: User study statistics.
Sub. Cons. Bg. Cons. LPIPS FVD
  Only fine-tuning 89.57 92.09 0.3290 540.47
Linear interpolation 90.77 92.75 0.3046 481.41
Pretrained optical flow 91.52 93.44 0.2871 454.29
Flow-based VFI 92.65 94.34 0.2716 401.49
(a) Choice of motion guidance.
Latent Feature Sub. Cons. Bg. Cons. LPIPS FVD
  89.57 92.09 0.3290 540.47
91.87 93.92 0.2796 437.85
92.34 93.74 0.2811 424.50
92.65 94.34 0.2716 401.49
(b) Different levels of guidance injection.
Sub. Cons. Bg. Cons. LPIPS FVD
  Transformers 92.35 94.01 0.2750 431.21
Convolution 92.44 94.09 0.2745 426.85
Linear 92.49 94.17 0.2739 422.97
Average 92.65 94.34 0.2716 401.49
(c) Different ways to merge guidance.
Sub. Cons. Bg. Cons. LPIPS FVD
  All 92.17 93.51 0.2792 451.31
Decoder-only 91.74 92.97 0.2942 471.52
Encoder-only 92.65 94.34 0.2716 401.49
(d) Position of feature-level injection.
Table 4: Ablation experiments. The colored background indicates our default setting.
Effectiveness of motion guidance.
As a key contribution of our work, we thoroughly validate the efficacy of the proposed motion guidance. In Fig. 5, (a) displays the results from the DynamiCrafter, which suffers from significant motion instability and content inconsistency. (b) presents outcomes warped by the intermediate flow from [41]. Although noticeable blurring is evident, the warped intermediate frames offer valuable approximations of both motion and appearance. (c) shows results obtained through direct guidance integration using our method in a train-free manner, demonstrating substantial improvements in motion stabilization. However, this direct integration, lacking training, leads to alterations in feature distribution that affect the color and details of generated frames. Finally, (d) showcases the results of the MoG, which both incorporates guidance and fine-tunes the spatial layers. Our method effectively leverages this guidance, producing realistic motion trajectories and maintaining consistent content. More comparisons can be found in the supplementary materials.
Figure 5: Ablation on motion guidance integration.
Choice of motion guidance.
To select the appropriate motion guidance, we evaluate various approaches in Sec. 4.1 and compare their results with those obtained without motion guidance (only fine-tuning), as shown in Tab. 4(a). As discussed in Sec. 4.1, linear interpolation offers only marginal improvements in video quality and fidelity, whereas the intermediate flow from flow-based VFI achieves superior performance due to its task-oriented training and explicit occlusion modeling.
Different levels of guidance injection.
We introduce motion guidance at both the latent and feature levels. To validate the effectiveness of each level, we compare the performance of models with motion guidance introduced at only one level, or without motion guidance, as in Tab. 4(d). The results demonstrate that the injection of either level all significantly enhances video quality and fidelity metrics. Specifically, latent-level injection is more beneficial for background consistency, while feature-level injection improves subject consistency. The best performance is achieved when both levels are employed, allowing the generative model to leverage motion guidance at different granularities to facilitate the generation of intermediate frames.
Design details of feature-level injection.
We investigate the details of feature-level injection from two perspectives: how and where to merge motion guidance. Regarding how to merge guidance, we also devise methods that first concatenate the warped intermediate features with the original features, followed by learnable modules such as Transformer blocks, convolutions, or linear layers. Surprisingly, as shown in Tab. 4(c), a simple averaging yielded the best performance, possibly due to the substantial data requirements for the learnable modules to achieve strong generalization. In terms of where to merge motion guidance, we experiment with three configurations: injecting into all blocks (All), exclusively into expansive blocks (Decoder-only), and solely into contracting blocks (Encoder-only). As illustrated in Tab. 4(b), the Encoder-only achieves the best performance, while Decoder-only results in the poorest performance. This discrepancy may stem from the fact that modifications to the decoder can disrupt the powerful video generation capabilities from pre-training.
6 Conclusion
In this work, we have presented a novel generative frame interpolation framework, MoG, which enhances the model’s motion awareness by explicitly incorporating additional motion guidance. We first reveal that the intermediate flow in flow-based VFI can serve as a suitable motion guidance. Subsequently, we propose a simple yet effective strategy to inject this guidance into the generative network at both the latent and feature levels, significantly reducing the challenge of inferring the motion of input frames. Through extensive experiments in both real-world and animated scenes, we demonstrate that MoG achieves substantial improvements in video quality and fidelity. User studies further indicate that MoG exhibits superior overall visual quality.
Appendix
Figure 6: Additional ablation experiments on motion guidance integration.
A. More Visual Comparisons
To further demonstrate the improvement of our method, we provide video comparisons of different methods on the provided website. We also provide additional qualitative comparisons in Fig. 8.
B. Effectiveness of Introducing Motion Guidance
As shown in Fig. 6, we provide more examples similar to Fig. 5 to demonstrate the effectiveness of introducing motion guidance. By explicitly incorporating a rough estimate of the motion between input frames, our method effectively stabilizes the generated motion and reduces the inconsistency in content between the generated video and the input frames.
C. Discussion on Ineffective Motion Guidance
Our method significantly enhances video generation quality primarily by leveraging the coarse intermediate frame estimates provided by motion guidance. To present a comprehensive analysis, it is essential to demonstrate our method’s performance in scenarios where motion guidance is ineffective. As illustrated in the two examples in Fig. 7, when the motion in the input frames is overly complex, the intermediate flow fails to deliver effective motion guidance. Notably, despite the ineffectiveness of motion guidance in these cases, our approach still yields high-quality results. This indicates that our method, MoG, does not rely entirely on motion guidance for generating identical motion. Instead, it dynamically integrates the effective information from motion guidance to assist in the inference of real motion. This aligns perfectly with our original intent.
D. Limitations and Future Work
Despite MoG has achieved non-trivial improvement in generation quality across various scenes, there still several limitations warrant further exploration. Firstly, our approach is built upon the U-Net architecture of the DynamiCrafter model. However, the video generation capabilities of DynamiCrafter has lagged behind recently DiT-based [22] video generation models [40, 4], which constrains our performance ceiling. Investigating our method within the new framework presents a promising avenue for future work. Secondly, our approach relies on a pre-trained flow-based VFI model, meaning that the quality of its outputs may impact the effectiveness of motion guidance. For instance, flow-based VFI struggles to establish correspondences for extremely large motion [15]. Enhancing the generalizability of flow-based VFI across diverse scenes will also benefit our method moving forward.
Figure 7: Instances of ineffective guidance.
Figure 8: Additional qualitative comparison on real-world and animation scenes.

### v2

@@
 Abstract
-Generative frame interpolation, empowered by large-scale pre-trained video generation models, has demonstrated remarkable advantages in complex scenes.
-However, existing methods heavily rely on the generative model to independently infer the correspondences between input frames—an ability that is inadequately developed during pre-training.
-In this work, we propose a novel framework, termed Motion-aware Generative frame interpolation (MoG), to significantly enhance the model’s motion awareness by integrating explicit motion guidance.
-Specifically we investigate two key questions: what can serve as an effective motion guidance, and how we can seamlessly embed this guidance into the generative model.
-For the first question, we reveal that the intermediate flow from flow-based interpolation models could efficiently provide task-oriented motion guidance.
-Regarding the second, we first obtain guidance-based representations of intermediate frames by warping input frames’ representations using guidance, and then integrate them into the model at both latent and feature levels.
-To demonstrate the versatility of our method, we train MoG on both real-world and animation datasets.
-Comprehensive evaluations show that our MoG significantly outperforms the existing methods in both domains, achieving superior video quality and improved fidelity.
-Figure 1: Examples of frame interpolation in real-world and animation scenes.
-Compared to other methods, our approach, MoG, exhibits superior stability in motion and consistency in content relative to the input frames.
-For optimal viewing, please zoom in.
+Flow-based frame interpolation methods ensure motion stability through estimated intermediate flow but often introduce severe artifacts in complex motion regions.
+Recent generative approaches, boosted by large-scale pre-trained video generation models, show promise in handling intricate scenes.
+However, they frequently produce unstable motion and content inconsistencies due to the absence of explicit motion trajectory constraints.
+To address these challenges, we propose Motion-aware Generative frame interpolation (MoG) that synergizes intermediate flow guidance with generative capacities to enhance interpolation fidelity.
+Our key insight is to simultaneously enforce motion smoothness through flow constraints while adaptively correcting flow estimation errors through generative refinement.
+Specifically, we first introduce a dual guidance injection that propagates condition information using intermediate flow at both latent and feature levels, aligning the generated motion with flow-derived motion trajectories.
+Meanwhile, we implemented two critical designs, encoder-only guidance injection and selective parameter fine-tuning, which enable dynamic artifact correction in the complex motion regions.
+Extensive experiments on both real-world and animation benchmarks demonstrate that MoG outperforms state-of-the-art methods in terms of video quality and visual fidelity.
+Our work bridges the gap between flow-based stability and generative flexibility, offering a versatile solution for frame interpolation across diverse scenarios.
+Figure 1: Examples of frame interpolation in real-world and animation scenes. Compared to other methods, our approach, MoG, exhibits superior stability in motion and consistency in appearance details.
 *Work is done during internship at Tencent PCG.  ‡Equal contribution.  †Corresponding author (lmwang@nju.edu.cn).
 1 Introduction
-Video Frame Interpolation (VFI), which seeks to synthesize intermediate frames between two input frames, has garnered significant attention in recent years due to its capacity for enhancing video frame rates in video post-processing.
-Flow-based VFI methods [10, 41, 42, 15, 13] predominantly rely on correspondence information through estimating the motion between input frames—termed intermediate flow—to warp information and generate intermediate frames.
-However, these approaches are primarily optimized for rigid motion scenarios.
-Consequently, when confronted with complex scenes, the interpolated frames often exhibit pronounced blurring and artifacts, as the results of the flow-based method EMA-VFI [41] shown in Fig.
+Video Frame Interpolation (VFI), which aims to synthesize intermediate frames between two input frames, has garnered significant attention in recent years due to its capacity for enhancing video frame rates in video post-processing.
+Flow-based VFI methods [9, 45, 46, 17, 14] predominantly rely on estimating the motion between input frames—termed intermediate flow—to warp condition information and generate intermediate frames.
+Conventionally, the generated intermediate frames follow the motion trajectories encoded in the estimated intermediate optical flow, yielding temporally coherent video sequences.
+However, in scenarios involving complex motions such as object deformations, accurate intermediate flow estimation becomes infeasible, leading to pronounced artifacts in corresponding regions.
+As demonstrated in Fig.
+1 by the flow-based method EMA-VFI [45], while the overall results exhibit temporal smoothness, complex regions manifest blurring and ghosting artifacts.
+Recent advancements [38, 39, 37, 4] have increasingly focused on leveraging video generation models [39, 2] for frame interpolation, capitalizing on their robust generative capacities in dynamic scenes.
+While notable improvements have been demonstrated in complex scenarios [38], current approaches rely exclusively on generative models to infer inter-frame correspondences—a capability that remains underdeveloped during generative pre-training.
+Consequently, these methods often produce unstable motion patterns and inconsistencies with input frames due to the lack of explicit motion trajectory constraints, as illustrated by DynamiCrafter [39] and ToonCrafter [38] in Fig.
 1.
-To overcome these limitations, recent advancements [34, 35, 33, 3] have shifted towards leveraging video generation models [35, 2] for frame interpolation, capitalizing on their strong generative capabilities in dynamic scenes.
-Despite the notable improvements exhibited in complex scenarios [34], current approaches still face a key challenge: they fail to explicitly exploit the dynamics between the two input frames.
-Instead, they solely rely on the generative model to infer the correspondences between input frames by itself—an ability that is insufficiently nurtured during the generative pre-training.
-Consequently, the videos generated by these methods often suffer from unstable motion and incoherent content to input frames, as shown by DynamiCrafter [35] and ToonCrafter [34] in Fig.
+In this work, we introduce a new framework, Motion-aware Generative frame interpolation (MoG), which integrates intermediate flow with generative capacities to to enhance interpolation fidelity.
+MoG is designed to to bridge the gap between flow-based stability and generative flexibility, offering a versatile solution for frame interpolation across diverse scenarios.
+Our core idea is to enforce motion smoothness via flow constraints while rectifying complex motion regions through generative refinement simultaneously.
+To attain this objective, we tackle two crucial questions: how to seamlessly incorporate flow constraints into the generative model, and how to endow the generative model with the ability to dynamically correct flow errors.
+For the first question, to ensure that the generated motion trajectories are guided by the estimated intermediate flow, we introduce dual guidance injection.
+At both the latent and feature levels, we warp the information of the input frames using the intermediate flow and propagate it into the generation process of intermediate frames, serving as an explicit motion guidance for inferring motion.
+Notably, compared with ControlNet-like designs [47, 50], our design requires no additional parameters and thereby better preserves the pre-trained capabilities.
+To address the second challenge, we introduce two critical designs: encoder-only guidance injection and selective parameter fine-tuning.
+Specifically, guidance is exclusively injected at the encoder stage of the generative model, enabling the decoder to adaptively adjust the generation process.
+Furthermore, we only fine-tune the spatial layers to adapt to guidance fusion while preserving motion modeling capabilities in temporal layers.
+To thoroughly evaluate the versatility of our approach, we adapt MoG for both real-world and animation scenes.
+Experimental results demonstrate that MoG substantially outperforms existing generative interpolation models in both domains.
+Specifically, the interpolated videos generated by MoG exhibit superior motion stability and improved content consistency, as visually demonstrated in Fig.
 1.
-In this work, we introduce a new framework, Motion-aware Generative frame interpolation (MoG), marking the first explicit incorporation of motion guidance between input frames to enhance the motion awareness of generative models.
-Our approach significantly alleviates the difficulty of inferring the dynamics between input frames to generate the realistic motion.
-To achieve this goal, we address two pivotal questions: what can serve as an effective motion guidance, and how to seamlessly and harmlessly integrate this guidance into the generative model.
-For the first question, we propose leveraging the intermediate flow from flow-based frame interpolation models [41] as explicit motion guidance.
-The intermediate flow can be utilized to obtain coarse approximations of intermediate frames under the assumption of rigid motion, which can serve as effective guidance to inspire and stabilize the inference of realistic motion in the generation process.
-Moreover, compared to direct estimation by pre-trained optical flow models [36, 28, 37], the intermediate flow is task-oriented trained [38], making it inherently more suitable for generating intermediate frames.
-Flow-based interpolation models also specifically account for occlusion masks [41] between the two input frames, enabling a more accurate aggregation of information into the intermediate frame—something that optical flow often fails to achieve.
-To address the second question, we propose a simple yet effective strategy: warping the information from input frames to derive the guidance-based intermediate frame representations, which are subsequently integrated into the pre-trained model as explicit motion cues.
-This integration is performed at both the latent and feature levels, enabling the model to be fully aware of the correspondences between input frames.
-Specifically, at the latent level, we concatenate a collection of auxiliary latents to the input at each denoising step [8], which includes the latents from the input frames as well as the warped intermediate latents.
-In terms of feature-level integration, we reuse the pre-trained temporal layers [35] to temporally smooth the warped features, ensuring their alignment within the original feature space, and then incorporate them into the network.
-Notably, our design requires no additional parameters and achieves improved performance with few fine-tuning steps.
-To comprehensively evaluate the versatility of our method, we develop MoG for both the real-world and animation scenes.
-Furthermore, we meticulously curate two testing datasets, VFIBench-Real and VFIBench-Ani, to assess the performance of each scenes.
-Experimental results indicate that MoG significantly outperforms the existing generative interpolation models in both qualitative and quantitative aspects.
-Qualitatively, the videos generated by MoG exhibit enhanced motion stability and greater content consistency, as evidenced in Fig.
-1.
-Quantitatively, MoG demonstrates the obvious improvements in video quality metrics and superior fidelity when compared to ground truth videos.
 Our contributions are summarized as follows:
-  • We introduce a new generative frame interpolation framework, MoG, which is the first to incorporate motion guidance to enhance the motion awareness.
-  • We demonstrate that the intermediate flow from flow-based VFI can effectively serve as a motion guidance, and we design a simple and efficient method to integrate this prior into the network at both the latent and feature levels.
-  • We validate the effectiveness of MoG in both real-world and animated scenarios, with experimental results showing that MoG significantly outperforms existing open-source generative frame interpolation methods.
+  • We propose a novel frame interpolation framework, MoG, which is the first to bridge the gap between flow-based stability and generative flexibility.
+  • We introduce dual-level guidance injection to constrain the generated motion with the motion trajectories derived from the flow.
+  • We implement encoder-only guidance injection and selective parameter fine-tuning to endow the generative model with the ability to dynamically correct flow errors.
+  • Experimental results showing that MoG significantly outperforms existing generative frame interpolation methods in both qualitative and quantitative aspects.
 2 Related Work
@@
 MoG consists of two parts.
-First, it extracts the motion guidance of the input frames using a pre-trained flow-based VFI model.
+First, it extracts the intermediate flow between input frames.
 Subsequently, this guidance is seamlessly injected into the generative model at both the latent and feature levels.
+Meanwhile, the generative model would adaptively rectifying the errors by two crucial designs, namely, encoder-only guidance injection and selective parameter fine-tuning.
 2.1 Flow-based Frame Interpolation
-Flow-based video interpolation, explicitly estimating the intermediate flow from the input frames to the intermediate frame, have become dominant in deterministic frame interpolation [17].
+Flow-based video interpolation, explicitly estimating the intermediate flow from the input frames to intermediate frames, have become dominant in deterministic frame interpolation [19].
 It can be broadly categorized into two classes based on how the intermediate optical flow is derived.
@@
 Similarly, EMA-VFI [45] enhanced the flow prediction by explicitly modeling the dynamics between frames through inter-frame cross-attention.
-However, both of classes are primarily designed under the assumption of rigid motion scenes.
-When confronted with complex motion scenarios, they often exhibit significant blurring and artifacts.
-In our work, we focus on the generative frame interpolation by leveraging the second class of flow-based video frame interpolation (EMA-VFI [41]) to provide explicit motion guidance.
+However, when confronted with complex motion scenarios, both of classes often exhibit significant blurring and ghosting artifacts.
+In this work, we leverage the intermediate flow to enhance the temporal smoothness of generative models.
+Meanwhile, we utilize generative models to rectify the errors of the intermediate flow in complex scenarios.
 2.2 Generative Frame Interpolation
 Recent work has begun to explore the use of large-scale pre-trained video generation models [2, 39], which excel at generating videos in complex dynamic scenes, for the VFI task.
-Current generative frame interpolation methods can be categorized into two types: the first [3, 33] employed pre-trained generative models to perform image-to-video tasks conditioned on the initial and final frames, subsequently merging the resulting videos to create the final interpolated frame.
+Current generative frame interpolation methods can be categorized into two types: the first [4, 37, 50, 43] employed pre-trained generative models to perform image-to-video tasks conditioned on the initial and final frames, subsequently merging the resulting videos to create the final interpolated frame.
 For example, GI [37] enhanced the motion stability by controlling the consistency of temporal correlations across the two generation processes.
-The second category [35, 34, 15, 32, 44] focused on fine-tuning video generation models specifically for interpolation, by integrating information from input frames into the model’s architecture and optimizing it for end-to-end interpolation.
+The second category [39, 38, 17, 35, 49, 3, 11, 31] focused on fine-tuning video generation models specifically for interpolation, by integrating information from input frames into the model’s architecture and optimizing it for end-to-end interpolation.
 DynamiCrafter [39] was trained for real-world interpolation, while ToonCrafter [38] was tailored for animated scenes.
 Although all these methods have demonstrated significant improvements in generating complex scenarios, they do not explicitly consider the correspondence between input frames, which complicates the motion inference of generative models.
-In contrast, we are the first to explicitly introduce the correspondence guidance to enhance the motion awareness of generative models and our method achieves superior video quality and fidelity.
+In contrast, we are the first to explicitly introduce the motion guidance to enhance the motion smoothness of generative models and our method achieves superior video quality and fidelity.
+Currently, there are also some works [42, 20, 16, 36] that use optical flow to assist generation in other tasks.
+They typically require the accurate optical flow between all adjacent frames to aid the generation process, whereas we only use the coarse intermediate flow between the two input frames to smooth the motion.
+Meanwhile, our method could dynamically correct the errors in the intermediate flow.
 3 Preliminaries
-Figure 3: Comparison of different methods for approximating intermediate frames.
-Linear interpolation merely overlays the input frames, while optical flow often introduces significant artifacts.
-In contrast, the frames warped through intermediate flow demonstrate superior stability, providing the generative model with robust correspondence information.
+Figure 3: Demonstration of MoG’s guidance correction capability.
+In both examples, DynamiCrafter or ToonCrafter struggle to generate temporally consistent motion in complex scenarios.
+While intermediate flow can provide valuable motion cues, it often introduces artifacts and fails to render fine appearance details.
+Leveraging the encoder-only injection design and selective parameter fine-tuning, MoG effectively integrates reliable motion information from intermediate flow while correcting its inaccuracies.
 3.1 Task Definition
@@
 After training, we can iteratively recover $\mathbf{\hat{z}}_{0}$ using the input conditions and pure noise $\mathbf{z}_{T}\sim\mathcal{N}(0,I)$, generating the video $\mathbf{\hat{x}}=\mathcal{D}(\mathbf{\hat{z}}_{0})$ via the decoder $\mathcal{D}$.
-The design of the denoising network $\epsilon_{\theta}$ follows an U-Net-like structure [26], consisting of contracting blocks, middle blocks, and expansive blocks.
+The design of the denoising network $\epsilon_{\theta}$ follows an U-Net-like structure [29], consisting of encoder blocks and decoder blocks.
 Each block comprises spatial and temporal layers.
@@
 4 Motion-Aware Generative VFI
-Currently, most generative interpolation methods [33, 3, 35, 34, 5] conduct frame interpolation by directly incorporating information from the input frames into the model’s input [5, 35] or decoder [35].
-However, in these methods, the correspondence between input frames, which is crucial for understanding and reasoning about the motion and appearance of the intermediate frames, have to be inferred independently by the generative model itself.
-This correspondence estimation capability has not been adequately cultivated in the pre-training of large-scale text-to-video generation or pure video generation tasks.
-Consequently, this insufficient exploration of correspondence leads to inconsistencies in motion and appearance of the generated videos relative to the input frames.
-To address this issue, we propose Motion-Aware Generative Frame Interpolation (MoG), the first approach to explicitly introduce correspondence information between the two frames into the generative model, thereby reducing the difficulty of the inference for motion and appearance for generative model.
-As illustrated in Fig.
-2, MoG comprises two parts: it first extracts an explicit motion guidance based on the input frames by a pre-trained flow-based VFI model [41]; it then seamlessly injects this guidance into the generative model at both the latent and feature levels.
-Notably, our method does not require any additional parameters; it merely necessitates minimal fine-tuning steps on spatial layers of the denoising network to compensate for changes in feature distribution.
-4.1 Choice of motion guidance
-As illustrated in Fig.
-3, in addition to the intermediate flow described in Eq.
-1, there are two other guidance for representing intermediate frames based on input frames.
-The first method is linear interpolation, which directly weights the representations of the initial and final frames according to the time step of the intermediate frame.
-While straightforward, this technique merely overlays the input frames, offering limited correspondence information.
-The second method employs a pre-trained optical flow estimator, such as [37], to derive the flow between input frames, which is then scaled to generate the intermediate flow.
-This approach often suffers from limited generalization due to being predominantly trained on synthetic datasets.
-Moreover, it cannot account for occlusion relationships between input frames.
-This can lead to artifacts in complex scenes, such as jagged edges and ghosting effects.
-In contrast, the intermediate flow defined in Eq.
-1 is specifically trained for the VFI task, making it inherently more suitable for representing the correlation between intermediate frames and the initial and final frames.
-It also predicts an occlusion mask $M^{n}$ for each frame.
-As illustrated in Fig.
-3, the intermediate flow allows for the approximation of intermediate frames, which, despite some blurriness, maintains stable motion trajectories and consistent object appearances.
-This stability alleviates the difficulties of real motion inference in generative models.
-This is further validated by our experiments in Sec.
-5.4.
-4.2 Motion Guidance Integration
-Due to the inability of diffusion models to directly use correspondence guidance, it is necessary to devise a strategy to inject these motion informaiton into the denoising network.
-To achieve this, we propose a simple yet effective approach that leverages the guidance, akin to the Eq.
-2, to coarsely estimate the representation of the intermediate frames from the representations of the start and end frames.
-This estimated representation is then seamlessly merged into the denoising network.
-Furthermore, to fully exploit the correspondence guidance across different granularities of the generation process, we conduct the guidance injection at both the latent and feature levels.
+4.1 Motivation
+Flow-based frame interpolation methods can generate temporally smooth videos with the assistance of intermediate flow.
+However, in regions with complex motions, accurate intermediate flow estimation remains challenging due to the lack of sufficient supervisory signals—the annotation of real-world optical flow is prohibitively expensive.
+This limitation leads to severe artifacts such as blurring and ghosting in flow-based approaches.
+In contrast, recent generative frame interpolation methods, benefiting from large-scale pre-training, demonstrate robust generative capabilities in complex scenarios.
+Unfortunately, they often suffer from incoherent motion and content inconsistencies with input frames.
+We attribute this to the fact that existing methods rely solely on generative models to infer motion trajectories between input frames—a capability insufficiently nurtured during generative pre-training.
+To bridge the gap between flow-based stability and generative flexibility, we introduce Motion-aware Generative frame interpolation (MoG), as illustrated in Fig.
+2.
+MoG employs the intermediate flow from flow-based methods as an explicitly motion guidance for generation, smoothing the motion of generated videos.
+Concurrently, MoG leverages the refinement capacities of generative models to automatically correct flow errors in complex motion regions.
+4.2 Dual Guidance Injection
+Since generative models cannot directly utilize the intermediate flow, it is imperative to devise a strategy for sufficiently injecting motion guidance into the denoising network.
+To this end, we propose dual guidance injection.
+Similar to the operation in Eq.
+2, we coarsely estimate the representation of the intermediate frames using the intermediate flow.
+Subsequently, the estimated representations are seamlessly integrated into the model at both the latent and feature levels.
 Latent-Level Injection.
-To introduce motion guidance at the latent level, we propose to coarsely estimate the latent code of intermediate frames by backward warping of the input frame’s latent code with the intermediate flow.
-Specifically, given the latent codes of the start and end frames $z_{0}^{1}$ and $z_{0}^{N}$, along with the motion guidance obtained via Eq.
-1, we estimate the latent code of the $n$-th intermediate frame:
+To incorporate flow constraints at the latent level, we propose to inject additional latent codes of intermediate frames into the input.
+We backward-warp the latent codes of the input frames, guided by the intermediate flow.
+Specifically, given the latent codes of the start and end frames $z_{0}^{1}$ and $z_{0}^{N}$, along with the motion guidance obtained through Eq.
+1, we estimate the latent code of the $n$-th intermediate frame as follows:
 $\bar{z}_{0}^{n}=\text{warp}(z_{0}^{1},f_{1\to n})\odot M^{n}+\text{warp}(z_{0}^{N},f_{N\to n})\odot(1-M^{n}).$ (6)
@@
 $\tilde{\mathbf{z}}_{t}=\left[\mathbf{z}_{t};\bar{\mathbf{z}}_{0}\right],\quad\tilde{\mathbf{z}}_{t}\in\mathbb{R}^{N\times(2\times C)\times h\times w}.$ (7)
-It is noteworthy that no additional parameters are required, as our base models, DynamiCrafter or ToonCrafter, have already designed to accommodate extra inputs; however, the latent codes of intermediate frames in $\mathbf{\bar{z}}_{0}$ are always zero at their methods.
+It is noteworthy that no additional parameters are required, as our base models, DynamiCrafter or ToonCrafter, have already designed to accommodate extra inputs; In contrast, in their methods, the latent codes of intermediate frames in $\mathbf{\bar{z}}_{0}$ are always set to zero.
 Feature-Level Injection.
-To effectively integrate the motion guidance in different granularities, we propose to inject guidance also in feature-level.
-Analogous to the latent-level, we estimate the features of intermediate frames based on the features $F^{0},F^{N}\in\mathbb{R}^{D\times H\times W}$ of the input frames:
+To effectively integrate the motion guidance cross different granularities, we propose to inject guidance also in feature-level.
+Analogous to the latent-level, we estimate the features of intermediate frames based on the features $F^{1},F^{N}\in\mathbb{R}^{D\times H\times W}$ of the input frames:
 $\bar{F}^{i}=\text{warp}(F^{1},f_{1\to i})\odot M^{n}+\text{warp}(F^{N},f_{N\to i})\odot(1-M^{n}).$ (8)
-In this equation, $\bar{F}^{i}$ represents the estimated features of the $i$-th frame under the correspondence prior.
+In this equation, $\bar{F}^{i}$ represents the estimated features of the $i$-th frame guided by the intermediate flow.
 Unfortunately, unlike in latent-level injection, direct concatenation of the warped intermediate features into the network is not feasible.
-To address this issue, we leverage the temporal layer of the denoising network, which enables us to smooth and align the estimated intermediate features with the original feature distribution without introducing additional parameters, as shown in Fig.
-2:
+To address this issue, we reuse the temporal layer of the denoising network, which enables us to smooth and align the estimated intermediate features $\mathbf{\bar{F}}$ with the original feature distribution without introducing additional parameters.
+As shown in Fig.
+2, the smoothed features $\mathbf{\hat{F}}$ is acquired by:
 […]
-Subsequently, we incorporate the smoothed features $\dot{F}$ into the original features:
-$\mathbf{\tilde{F}}=(\mathbf{F}+\mathbf{\hat{F}})/2.$ (10)
-Remarkably, our exploration (refer to Sec. 5.4) reveals that this simple averaging already allows the generative model to effectively utilize the introduced motion guidance.
+Subsequently, we incorporate the flow-derived features $\dot{F}$ into the original features:
+$\mathbf{\tilde{F}}=\phi(\mathbf{F},\mathbf{\hat{F}}).$ (10)
+Remarkably, our exploration (refer to Sec. 5.4) reveals that a simple averaging already allows the generative model to effectively utilize the introduced motion guidance.
+4.3 Guidance Correction
+As mentioned in Sec.
+4.1, the intermediate flow cannot be accurately predicted in regions with complex motions.
+To alleviate the potential adverse effects, we have implemented two crucial designs to endow the generative model with the ability to automatically rectify error-prone regions.
+Firstly, we adopt the design of encoder-only guidance injection.
+As depicted in Fig.
+2, we introduce feature-level guidance injection exclusively in the encoder blocks of the denoising network.
+In this way, the decoder can appropriately adjust and rectify the information from the encoder, thereby capitalizing more effectively on the valuable information embedded in the flow guidance.
+Secondly, we conduct selective parameter fine-tuning.
+We only fine-tune the spatial layers of all blocks.
+Given the inherent inaccuracy of the intermediate flow during the training phase, the model can learn how to dynamically harness the useful information within the guidance while concurrently acquiring the ability to repair flawed regions.
+Freezing the temporal layers serves to preserve the motion inference capability during pre-training and prevent the performance degradation associated with the feature-level guidance injection.
+Discussion.
+To validate the effectiveness of our proposed method, as shown in Fig.
+3, we showcase two examples for a comparison among the results derived from the intermediate flow, those of the original generative model, and the results produced by MoG.
+Evidently, MoG can fully exploit the motion cues encapsulated within the intermediate flow to generate videos with better temporal smoothness.
+Meanwhile, in regions with complex motions, MoG rectifies the inaccuracies from the intermediate flow and leverages the generative capabilities to yield more reasonable appearance details.
+More quantitative comparisons can be found in Sec.
+5.4.
 5 Experiment
 […]
-[…]
-[…]
-[…]
-[…]
-[…]
-Average
 Real Anime Real Anime Real Anime Real Anime Real Anime Real Anime Real Anime
       Flow-based VFI models
-RIFE [10] 86.63 86.92 91.67 94.25 97.92 98.46 99.28 99.38 45.18 48.22 46.15 48.11 77.81 79.22
-EMA-VFI [41] 87.85 88.97 91.91 95.12 97.39 98.29 99.11 99.29 46.68 50.20 47.96 52.24 78.48 80.69
+RIFE [9] 18.21 20.33 0.5672 0.7587 0.3601 0.3407 55.27 62.35 0.8304 0.8693 742.29 628.55 77.57 78.59
+EMA-VFI [45] 18.17 20.49 0.5731 0.7531 0.3619 0.3701 51.09 53.27 0.8411 0.8907 717.58 517.60 78.15 80.02
       Generative VFI models
-GI [33] 91.81 91.28 91.70 94.16 93.86 96.43 97.51 98.00 47.44 51.13 57.51 60.84 79.97 81.97
-TRF [3] 89.25 89.25 92.26 94.01 93.48 95.77 96.77 97.18 47.94 50.12 54.38 58.40 79.01 80.79
-DynamiCrafter [35] 89.06 91.94 92.97 96.39 48.47 58.24 79.51
-ToonCrafter [34] 91.78 95.42 95.63 96.75 51.44 64.37 82.57
-MoG (ours) 92.65 92.91 94.34 96.11 94.69 96.58 97.90 96.94 49.84 52.43 59.23 64.89 81.44 83.31
-Table 1: Quantitative comparison on video quality. Bold text indicates the best results in generative frame interpolation.
-[…]
-Real Anime Real Anime Real Anime Real Anime Real Anime Real Anime
-      Flow-based VFI models
-RIFE [10] 18.13 20.27 0.6094 0.7817 0.3745 0.3407 898.38 647.49 65.04 66.16 0.8228 0.8643
-EMA-VFI [41] 18.11 20.45 0.6118 0.7843 0.3763 0.3701 808.54 544.48 60.28 55.84 0.8363 0.8895
-      Generative VFI models
-GI [33] 15.95 18.04 0.5271 0.6971 0.3384 0.2891 521.00 449.31 36.06 46.18 0.8703 0.8710
-TRF [3] 15.43 16.49 0.5132 0.6744 0.3920 0.3470 624.63 481.02 42.48 53.95 0.8491 0.8731
-DynamiCrafter [35] 16.05 0.5225 0.3380 562.34 42.16 0.8634
-ToonCrafter [34] 18.01 0.7182 0.2944 425.71 40.63 0.9203
-MoG (ours) 17.82 19.44 0.5898 0.7434 0.2716 0.2615 401.49 351.41 31.26 33.73 0.9083 0.9320
-Table 2: Quantitative comparison on generation fidelity to the ground truth.
+LDMVFI [3] 17.17 18.39 0.5953 0.7175 0.3081 0.2860 41.47 46.18 0.8703 0.8710 479.63 435.17 78.91 80.21
+GI [37] 15.95 18.04 0.5271 0.6971 0.3384 0.2891 36.06 46.18 0.8703 0.8710 521.00 449.31 79.97 81.97
+TRF [4] 15.43 16.49 0.5132 0.6744 0.3920 0.3470 42.48 53.95 0.8491 0.8731 624.63 481.02 79.01 80.79
+DynamiCrafter [39] 16.05 0.5225 0.3380 42.16 0.8634 562.34 79.51
+ToonCrafter [38] 18.01 0.7182 0.2944 40.63 0.9203 425.71 82.57
+MoG (ours) 17.82 19.44 0.5898 0.7434 0.2716 0.2615 31.26 33.73 0.9083 0.9320 401.49 351.41 81.44 83.31
+Table 1: Quantitative comparison on VFIBench.
 5.1 Implementation Details
@@
 A well-interpolated video should not only be of high quality inherently but also maintain fidelity to the ground truth.
-For the former, we utilize six metrics from VBench [11]: subject consistency, background consistency, temporal flickering, motion smoothness, aesthetic quality, and imaging quality.
-These metrics collectively assess the intrinsic quality of the video.
+For the former, we adopt six metrics from VBench [10]: subject consistency, background consistency, temporal flickering, motion smoothness, aesthetic quality, and imaging quality.
+The average performance across all metrics is reported as VBench in Tab.
+1.
+These metrics collectively assess the intrinsic quality.
 For the latter, we employ six widely adopted metrics: PSNR, SSIM, LPIPS [48], FID [7], and the CLIP similarity score [27] for image-level comparison, and FVD [32, 33] for video-level comparison.
 Figure 4: Visual comparison on real-world and animation scenes.
-5.3 Comparative Analysis
-We evaluate our MoG by benchmarking it against state-of-the-art methods across two categories: flow-based interpolation methods, specifically RIFE [10] and EMA-VFI [41], and generative interpolation methods, including GI [33], TRF [3], DynamiCrafter [35] and ToonCrafter [34].
-Note that DynamiCrafter and ToonCrafter are tailored for real-world and cartoon animation data, respectively, and their performance is reported separately for each data type.
+5.3 Main Results and Analysis
+We evaluate our MoG by benchmarking it against state-of-the-art methods across two categories: flow-based interpolation methods, specifically RIFE [9] and EMA-VFI [45], and generative interpolation methods, including LDMVFI [3], GI [37], TRF [4], DynamiCrafter [39] and ToonCrafter [38].
+To conduct a more equitable comparison, we retrain the flow-based methods and LDMVFI [3] on our dataset.
+Note that DynamiCrafter and ToonCrafter are tailored for real-world and animation, respectively.
+Hence their performance is reported separately for each domain.
 Quantitative results.
 As shown in Tab.
-1 and Tab.
-2, compared to others generative VFI methods, MoG exhibits significant improvements in video quality and fidelity to ground truth, particularly in consistency-related metrics.
-This demonstrates that the introduced motion guidance indeed enables the generative model to better understand the correspondence between input frames.
-Compared to flow-based VFI models, our approach also demonstrates notable enhancements across most metrics; however, it lags in PSNR, SSIM, Temporal Flickering, and Motion Smoothness.
+1, compared to others generative VFI methods, MoG exhibits significant improvements in video quality and fidelity to ground truth.
+This indicates that our method can effectively utilize the intermediate flow to generate smooth motion, and meanwhile, it has successfully reduced the erroneous information within it.
+Compared to flow-based VFI models, our approach also demonstrates notable enhancements across most metrics; however, it lags in PSNR.
 We argue that this discrepancy arises because flow-based VFI often produces blurry results in complex motion scenarios (as illustrated in Fig.
 1), which can inflate these metrics while compromising actual visual quality [48].
-Qualitative results.
-We present qualitative comparisons with three generative VFI methods in Fig.
-4.
-Lacking explicit motion guidance, these methods struggle to accurately infer and understanding the correspondences between input frames, resulting in inconsistent content and unstable motion.
-In contrast, MoG achieves superior motion and visual quality in complex scenarios.
-More comparisons are available on the provided website in supplementary materials.
-User study.
-To further verify the advantage of our method, we also conduct a comprehensive user study.
-Participants are instructed to select the best-generated videos based on motion quality, temporal coherence, frame fidelity, and overall quality.
-We collect results from 27 participants and report the findings in Tab.
-3.
-Thanks to the explicit motion guidance, the study shows a clear preference for our method in all aspects.
-5.4 Ablation Study
-For brevity, we only conduct ablation experiments in real-world scenarios.
-Our analysis primarily relies on four metrics to evaluate different strategies: two pertaining to video quality, namely Subject Consistency and Background Consistency (abbreviated as Sub.
-Cons.
-and Bg.
-Cons.
-in Tab.
-4), and two metrics assessing fidelity between the video and ground truth, specifically LPIPS and FVD.
+We also conduct a comparison on the publicly available dataset Davis-7 [12].
+As shown in Tab.
+2, MoG achieves best performance across all metrics.
+  Davis-7 SSIM ($\uparrow$) LPIPS ($\downarrow$) FID ($\downarrow$)
+  LDMVFI [3] 0.4175 0.2765 56.28
+VIDIM [12] 0.4221 0.2986 53.38
+DynamiCrafter [39] 0.4785 0.3752 75.06
+MoG (ours) 0.5978 0.2641 51.94
+Table 2: Comparsion on Davis-7 [12].
 […]
@@
 Table 3: User study statistics.
+  GI [37] TRF [4] DynamiCrafter [39] MoG (ours)
+  Runtime (s) 385.19 141.41 33.42 34.08
+VBench ($\uparrow$) 79.97 79.01 79.51 81.44
+Table 4: Comparsion on computational efficiency.
+Qualitative results.
+We present qualitative comparisons with three generative VFI methods in Fig.
+4.
+Lacking explicit motion guidance, these methods struggle to accurately infer and understanding the correspondences between input frames, resulting in inconsistent content and unstable motion.
+In contrast, MoG achieves superior motion and visual quality in complex scenarios.
+More comparisons are available in supplementary materials.
+User study.
+To further verify the advantage of our method, we also conduct a comprehensive user study.
+Participants are instructed to select the best-generated videos based on motion quality, temporal coherence, frame fidelity, and overall quality.
+We collect results from 27 participants and report the findings in Tab.
+3.
+Thanks to the explicit motion guidance, the study shows a clear preference for our method in all aspects.
+Computational efficiency .
+We further compare the computational efficiency of different generative frame interpolation methods.
+As shown in Tab.
+4, the testing resolution is unified as $16\times 512\times 320$.
+Compared with TRF and GI, MoG not only exhibits superior performance but also significantly improves computational efficiency.
+When compared with our baseline model, DynamicCrafter, MoG have substantially enhanced the quality of the generated videos while only slightly increasing the computational load.
+5.4 Ablation Study
+For brevity, we only conduct ablation experiments in real-world scenarios.
+Our analysis primarily relies on four metrics to evaluate different strategies: two pertaining to video quality, namely Subject Consistency and Background Consistency (abbreviated as Sub.
+Cons.
+and Bg.
+Cons.
+in Tab.
+5), and two metrics assessing fidelity between the video and ground truth, specifically LPIPS and FVD.
 Sub. Cons. Bg. Cons. LPIPS FVD
@@
 Table 5: Ablation experiments. The colored background indicates our default setting.
-Effectiveness of motion guidance.
-As a key contribution of our work, we thoroughly validate the efficacy of the proposed motion guidance.
-In Fig.
-5, (a) displays the results from the DynamiCrafter, which suffers from significant motion instability and content inconsistency.
-(b) presents outcomes warped by the intermediate flow from [41].
-Although noticeable blurring is evident, the warped intermediate frames offer valuable approximations of both motion and appearance.
-(c) shows results obtained through direct guidance integration using our method in a train-free manner, demonstrating substantial improvements in motion stabilization.
-However, this direct integration, lacking training, leads to alterations in feature distribution that affect the color and details of generated frames.
-Finally, (d) showcases the results of the MoG, which both incorporates guidance and fine-tunes the spatial layers.
-Our method effectively leverages this guidance, producing realistic motion trajectories and maintaining consistent content.
-More comparisons can be found in the supplementary materials.
-Figure 5: Ablation on motion guidance integration.
-Choice of motion guidance.
-To select the appropriate motion guidance, we evaluate various approaches in Sec.
-4.1 and compare their results with those obtained without motion guidance (only fine-tuning), as shown in Tab.
-4(a).
-As discussed in Sec.
-4.1, linear interpolation offers only marginal improvements in video quality and fidelity, whereas the intermediate flow from flow-based VFI achieves superior performance due to its task-oriented training and explicit occlusion modeling.
-Different levels of guidance injection.
+Figure 5: An example of manually altering intermediate flow.
+Effectiveness of intermediate flows.
+To validate that intermediate flows can indeed impose motion constraints on generative models, we attempt to manually modify the intermediate flows to control the motion.
+As depicted in Fig.
+5, given the same start and end frames, MoG would generate a nearly static video by default.
+When we manually alter the intermediate flows of the ball to move vertically, the generated frames change accordingly.
+In addition, we explore two other potential forms of motion guidance.
+One is the linear interpolation of the input frames, and the other is the flow from a pre-trained optical flow estimator [40].
+As shown in Fig.
+6, the intermediate flows exhibit smoother motion and fewer errors.
+We attribute this to the fact that intermediate flows are task-oriented flows [41], which are more suitable for the frame interpolation.
+Moreover, the predicted occlusion masks in Eq.
+1 can also enhance the accuracy of the guidance.
+The performance advantages presented in Tab.
+5(a) further verify this conclusion.
+Figure 6: Visualizations on different methods as guidance.
+Dual guidance injection.
 We introduce motion guidance at both the latent and feature levels.
-To validate the effectiveness of each level, we compare the performance of models with motion guidance introduced at only one level, or without motion guidance, as in Tab.
-4(d).
+To validate the effectiveness of each level, we compare the performance with motion guidance introduced at only one level, or without motion guidance, as in Tab.
+5(d).
 The results demonstrate that the injection of either level all significantly enhances video quality and fidelity metrics.
 Specifically, latent-level injection is more beneficial for background consistency, while feature-level injection improves subject consistency.
-The best performance is achieved when both levels are employed, allowing the generative model to leverage motion guidance at different granularities to facilitate the generation of intermediate frames.
-Design details of feature-level injection.
-We investigate the details of feature-level injection from two perspectives: how and where to merge motion guidance.
-Regarding how to merge guidance, we also devise methods that first concatenate the warped intermediate features with the original features, followed by learnable modules such as Transformer blocks, convolutions, or linear layers.
+The best performance is achieved when both levels are employed, allowing the generative model to leverage motion guidance at different granularities.
+Regarding the design of feature-level injection, we also attempt methods that first concatenate the warped intermediate features with the original features, followed by learnable modules such as Transformer blocks [34], convolutions, or linear layers.
 Surprisingly, as shown in Tab.
-4(c), a simple averaging yielded the best performance, possibly due to the substantial data requirements for the learnable modules to achieve strong generalization.
-In terms of where to merge motion guidance, we experiment with three configurations: injecting into all blocks (All), exclusively into expansive blocks (Decoder-only), and solely into contracting blocks (Encoder-only).
-As illustrated in Tab.
-4(b), the Encoder-only achieves the best performance, while Decoder-only results in the poorest performance.
-This discrepancy may stem from the fact that modifications to the decoder can disrupt the powerful video generation capabilities from pre-training.
+5(c), a simple averaging yielded the best performance, possibly due to the substantial data requirements for the learnable modules to achieve strong generalization.
+Guidance correction.
+We conduct comprehensive ablation experiments on designs proposed in Sec.
+4.3.
+First, regarding the integration location of motion guidance, we evaluate three configurations: guidance injection into all blocks (All), exclusive injection into decoder blocks (Decoder-only), and sole injection into encoder blocks (Encoder-only).
+As demonstrated in Tab.
+5(b), Encoder-only achieves the optimal performance, while Decoder-only yields the lowest performance.
+This discrepancy aligns with our hypothesis that maintaining decoder guidance-free allows the model to effectively rectify flow-constrained information from the encoder, ensuring the generated videos adhere more closely to pre-trained distributions.
+Furthermore, as illustrated in Fig.
+7, we validate the effectiveness of selective parameter fine-tuning.
+Results show that while intermediate flow guidance can improve temporal smoothness, it still leads to artifacts in complex regions.
+Selective fine-tuning substantially enhances the generated video quality by enabling dynamic guidance adaptation in spatial layers, while preserving the motion generation capabilities through frozen temporal layers.
+Figure 7: Visualizations on selective parameter fine-tuning.
 6 Conclusion
-In this work, we have presented a novel generative frame interpolation framework, MoG, which enhances the model’s motion awareness by explicitly incorporating additional motion guidance.
-We first reveal that the intermediate flow in flow-based VFI can serve as a suitable motion guidance.
-Subsequently, we propose a simple yet effective strategy to inject this guidance into the generative network at both the latent and feature levels, significantly reducing the challenge of inferring the motion of input frames.
+In this work, we present a novel generative frame interpolation framework, MoG, which simultaneously enhances the motion smoothness by intermediate flows and adaptively correcting flow errors through generative refinement.
+We first propose dual guidance injection to introduce flow constraints into the generative models at both the latent and feature levels.
+Then we conduct encoder-only guidance injection and selective parameters fine-tuning for guidance correction.
 Through extensive experiments in both real-world and animated scenes, we demonstrate that MoG achieves substantial improvements in video quality and fidelity.
-User studies further indicate that MoG exhibits superior overall visual quality.
 Appendix
-Figure 6: Additional ablation experiments on motion guidance integration.
 A. More Visual Comparisons
-To further demonstrate the improvement of our method, we provide video comparisons of different methods on the provided website. We also provide additional qualitative comparisons in Fig. 8.
-B. Effectiveness of Introducing Motion Guidance
-As shown in Fig.
-6, we provide more examples similar to Fig.
-5 to demonstrate the effectiveness of introducing motion guidance.
-By explicitly incorporating a rough estimate of the motion between input frames, our method effectively stabilizes the generated motion and reduces the inconsistency in content between the generated video and the input frames.
-C. Discussion on Ineffective Motion Guidance
-Our method significantly enhances video generation quality primarily by leveraging the coarse intermediate frame estimates provided by motion guidance.
-To present a comprehensive analysis, it is essential to demonstrate our method’s performance in scenarios where motion guidance is ineffective.
-As illustrated in the two examples in Fig.
-7, when the motion in the input frames is overly complex, the intermediate flow fails to deliver effective motion guidance.
-Notably, despite the ineffectiveness of motion guidance in these cases, our approach still yields high-quality results.
-This indicates that our method, MoG, does not rely entirely on motion guidance for generating identical motion.
-Instead, it dynamically integrates the effective information from motion guidance to assist in the inference of real motion.
-This aligns perfectly with our original intent.
-D. Limitations and Future Work
+To further demonstrate the improvement of our method, we also provide additional qualitative comparisons in Fig. 8.
+B. Limitations and Future Work
 Despite MoG has achieved non-trivial improvement in generation quality across various scenes, there still several limitations warrant further exploration.
@@
 However, the video generation capabilities of DynamiCrafter has lagged behind recently DiT-based [25] video generation models [44, 5], which constrains our performance ceiling.
-Investigating our method within the new framework presents a promising avenue for future work.
-Secondly, our approach relies on a pre-trained flow-based VFI model, meaning that the quality of its outputs may impact the effectiveness of motion guidance.
-For instance, flow-based VFI struggles to establish correspondences for extremely large motion [15].
+Unfortunately, we currently lack the necessary resources and data to work with the latest models.
+Investigating our method within the new models presents a promising direction for future work.
+Secondly, MoG relies on the flow-based VFI model, meaning that the quality of its outputs may impact the effectiveness of motion guidance.
 Enhancing the generalizability of flow-based VFI across diverse scenes will also benefit our method moving forward.
-Figure 7: Instances of ineffective guidance.
 Figure 8: Additional qualitative comparison on real-world and animation scenes.